### v1

# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

Abstract
Deep learning has driven significant progress in object detection using Synthetic Aperture Radar (SAR) imagery. Existing methods, while achieving promising results, often struggle to effectively integrate local and global information, particularly direction-aware features. This paper proposes SAR-Net, a novel framework specifically designed for global fusion of direction-aware information in SAR object detection. SAR-Net leverages two key innovations: the Unity Compensation Mechanism (UCM) and the Direction-aware Attention Module (DAM). UCM facilitates the establishment of complementary relationships among features across different scales, enabling efficient global information fusion. Among them, Multi-scale Alignment Module (MAM) and distinct Multi-level Fusion Module (MFM) enhance feature integration by capturing both texture detail and semantic information. Then, Multi-feature Embedding Module (MEM) feeds back global features into the primary branches, further improving information transmission. Additionally, DAM, through bidirectional attention polymerization, captures direction-aware information, effectively eliminating background interference. Extensive experiments demonstrate the effectiveness of SAR-Net, achieving state-of-the-art results on aircraft (SAR-AIRcraft-1.0) and ship datasets (SSDD, HRSID), confirming its generalization capability and robustness.
Index Terms:
Deep learning, SAR object detection, global fusion, direction-aware attention.
I Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring[1, 2]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships. In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to designing object detection methods using Convolutional Neural Networks (CNNs), yielding notable results [3, 4, 5, 6, 7, 8, 9]. However, the unique challenges posed by SAR images, characterized by their multi-scale and rotational nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands at a critical challenge in this field. Recent advancements in SAR image object detection have prominently featured deep learning approaches. Addressing the challenges of detecting objects in varying orientations, Li et al. [10] developed the Spatial Frequency Feature Fusion Network. This innovative network utilizes a Feature Pyramid Network (FPN) [11] to create spatial multi-scale features, specifically for ship objects, through a top-down hierarchical approach. Additionally, it employs a polar coordinate Fourier transform to attain rotation-invariant features in the frequency domain. In another notable work, Zhang et al. [12] introduced the Quad-Feature Pyramid Network (Quad-FPN), an architectural innovation consisting of four feature pyramid networks. This design effectively confronts issues such as background noise, multi-scale object representation, and the subtle features of small boats. Complementing these advancements, Li et al. [13] developed the Adjacent Feature Fusion (AFF) module, which strategically integrates local shallow features into adjacent layers. This selective and adaptive approach significantly enhances feature fusion, leading to improved representation of objects in SAR images.
Figure 1: Comparison of information fusion and transmission between our method and existing methods. In the figure, U represents unity operation, C represents feature compensation, and F represents feature fusion. Our perspective enables each layer to focus on global information, while existing fusion methods only focus on local information, resulting in information loss.
Current research in SAR image object detection primarily focuses on utilizing a FPN or its variants for feature fusion, offering a partial solution to multi-scale challenges. However, a significant limitation of these methods is their lack of global information integration at each layer. As illustrated in Fig. 1 (right side), local information flow between adjacent layers leads to a dilution of information across non-adjacent layers, resulting in information loss [14]. Recognizing the importance of global information, Bai et al. [15] proposed a Global Context-Guided Feature Balance Pyramid (GC-FBP), incorporating a global context module to bridge distant relationships between ship objects and their background. Yet, their approach of simply stacking context information onto the original feature map impedes effective interplay of information. Our method, depicted on the left side of Fig. 1, diverges from traditional algorithms that primarily depend on straightforward concatenation of adjacent feature layers. Instead, we adopt a global information fusion strategy, allowing for feature integration across various scales. This approach not only enhances the effectiveness and efficiency of the fusion process but also minimally impacts information loss during transmission. Traditional SAR object detectors, which predominantly use horizontal boxes, typically capture features of directional objects through an attention module [16] or by transforming features into the frequency domain [10]. However, these methods tend to neglect the correlation between the object and its contextual environment. This oversight often leads to ambiguous differentiation between the object and the background, with the resultant features being significantly noisy. Furthermore, these techniques generally overlook the crucial role of positional information. In contrast, our approach consider not only the object’s features but also the surrounding context. By integrating directional and positional information and embedding channel attention, our approach offers a comprehensive solution. We have distilled the problem of SAR image object detection into a broader challenge: addressing the issues of multi-scale and directional objects in predominantly background-dominated images. We base our considerations on three key premises: (1) the significant variation in object scales in SAR images. (2) the need to consider both shallow textural and deep semantic information of objects. (3) the directional nature of objects coupled with background interference. To address these challenges, we introduce a unity compensation mechanism, which includes three components: the Multi-scale Alignment Module (MAM), the Multi-level Fusion Module (MFM), and the Multi-feature Embedding Module (MEM). MAM initially unites multi-scale features across different spatial hierarchical levels. MFM then globally merges high-level semantic information with low-level detailed information, providing a comprehensive overview of the united features. Lastly, MEM compensates for any original feature to enrich their representational information. This multifaceted approach not only improves the detection of multi-scale objects but also significantly enhances information fusion capabilities. In addressing the third premise, we present a direction-aware attention module equipped with bidirectional attention polymerization capabilities. This innovative module employs deformable convolution and global pooling operations to extract directional and positional information along both vertical and horizontal axes of the input features. It then generates a direction-aware attention map, which is effectively integrated with the original input features. This approach adeptly tackles the challenge of accurately discerning directional and positional information amidst intricate background interference. Our contributions can be summarized in three primary points: $\bullet$ We introduce SAR-Net, a novel framework specifically designed for SAR image object detection. SAR-Net incorporates a unity compensation mechanism that globally unites multi-scale features of objects while compensating for the original feature. This ensures that global information is inherently integrated within features of different scales. $\bullet$ We design a direction-aware attention module with bidirectional attention polymerization capabilities. This module extracts object features while obtaining directional and positional information, enabling the identification of object rotation poses. Such a design is particularly effective in complex SAR images, where the directional attributes of objects are often obscured by significant background interference. $\bullet$ To validate the effectiveness, generalizability, and robustness of SAR-Net, we conduct comprehensive comparative and ablation experiments using three publicly available benchmark datasets for SAR aircraft and ship detection. The results of these experiments clearly show that our SAR-Net framework outperforms other state-of-the-art methods in all three datasets. The remainder of this paper is organized as follows. Section II reviews the development of object detection in the field of SAR images and related literature on solving multi-scale problems. Section III details our proposed framework and modules. In Section IV, the experimental setup is described, and qualitative and quantitative results are reported. Finally, conclusions are drawn in Section V.
II Related Work
II-A SAR Object Detection
Over the past few years, object detection in SAR images has aroused rising attention. Traditional detection algorithms, adhering to Constant False Alarm Rate (CFAR) principles [17], operate on a statistical model, dynamically setting the detection threshold based on a specified false alarm rate and the statistical distribution of background clutter. Consequently, a multitude of CFAR-based algorithms have surfaced. In a recent work [18], a novel decomposition approach was introduced to scrutinize the scattering between ships and the sea background. This decomposition technique, when integrated with CFAR, proves effective in ship detection. In a different vein, Lang et al. [19] enhanced the representation of pixels by incorporating spatial and intensity information, markedly improving the distinguishability between ships and background elements. However, the advent of high-resolution SAR images has introduced challenges such as complex ground interference and textured scenes accompanying multi-scale objects. This complexity renders the calculation of significant background clutter pixels in large-scale SAR images time-consuming. In response to the challenges of multi-scale object detection in SAR images, scholars have proposed several refined CFAR object detection algorithms. Examples include the scale sliding window object detection algorithm [20], the optimized super-pixel CFAR object detection algorithm [21], and the iterative CFAR object detection algorithm [22]. For detecting objects at larger scales, Zhai et al. [23] devised a ship detection algorithm emphasizing saliency and contextual information processing. This approach excels in focusing on sizable ships and background objects with prominent features, yielding commendable results. Despite these achievements, traditional methods still exhibit limitations in detecting small objects, necessitate manual parameter configuration, and heavily rely on manual design, presenting challenges for robustness and hindering their applicability to other SAR datasets. Consequently, there is a pressing need for more adaptive and automated approaches in SAR object detection. Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies. Shifting focus to SAR ship detection and identification, Zhang et al. [24] proposed a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of SAR aircraft detection and recognition, Zhao et al. [25] introduced an attention module, strategically fusing refined low-level texture features with high-level semantic features. This fusion strategy enhances aircraft detection rates by leveraging complementary information. Furthermore, Zhao et al. [26] devised a multi-branch dilated convolution feature pyramid method. This innovative approach minimizes redundancy while accentuating critical aircraft features through the establishment of dense connections. In a related endeavor, Zhao et al. [27] put forth a multitask learning framework for object detection (MLDet) tailored for ship detection in SAR images. This end-to-end framework encompasses tasks such as object detection, speckle suppression, and object segmentation. The collaborative execution of these tasks contributes to an enhanced detection performance. Addressing the specific challenge of large-scale SAR images, Yang et al. [28] concentrated on a multi-channel aggregation framework. This framework holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition. Although progress has been made, scale variation across object instances is still one of the most knotty problems in SAR object detection.
Figure 2: Overview of the proposed SAR-Net framework, here, the DAVgg Block embeds our designed direction-aware attention module, and the DA Block consists of DAVgg Blocks, combined to extract effective direction-aware information. $Ci$, $Ki$, and $Bi$ represent the multi-scale features at different stages.
II-B Multi-scale Object Detection
Within the domain of deep learning, the application of image pyramids initially emerged as a solution to address multi-scale challenges in object detection, driven by its intuitive appeal [29, 30, 31, 32, 33]. Approaches such as SNIP [34] and SNIPER [35] stand out for their selection of specific scales for each resolution during multi-scale training. However, it’s noteworthy that the utilization of image pyramids in these approaches substantially escalates inference time, presenting a trade-off between accuracy and efficiency. An alternative strategy seeks to incorporate in-network feature pyramids, offering an approximation of image pyramids with reduced computational costs. Notably, the Single-Shot Detector (SSD) [36] exemplifies this concept by leveraging multi-scale feature maps from different layers. This enables the detection of objects spanning various scales at each feature layer. Despite the computational advantages, it is essential to acknowledge that bottom-up sampling methods, like those employed in SSD, may exhibit diminished accuracy when confronted with small objects. This limitation arises from the inherent inadequacy of semantic information in low-level features, particularly impacting the precision of small object detection. Balancing the trade-offs between computational efficiency and accuracy remains a critical consideration in the ongoing evolution of multi-scale object detection methodologies within the realm of deep learning. To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [11] was introduced. FPN employs a bottom-up pathway, a top-down pathway, and lateral connections to effectively merge features of diverse resolutions and scales. Building upon FPN, the Path Aggregation Network (PANet) [37] introduced a bottom-up path to enhance information fusion between different levels. Researchers at FSAF [38] identified a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF extracts the feature representation of the region of interest (RoI) at a single feature level, overlooking any valuable information present on other feature levels of the pyramid. The SA-FPN [39], amalgamating Top-Down style FPN and Bottom-Up style FPN, combines the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrated information at all levels through common operations like element-wise max or sum, these fusion mechanisms did not consider semantic differences among feature layers. EfficientDet [40] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. YOLO-ASFF [41] employed an adaptive spatial feature fusion technique, learning weights during training to filter out inconsistencies. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [42] provided insights into this structure. Moreover, [43] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [44] proposed a parallel FPN structure for object detection with bi-directional fusion, while [45] introduced a refined FPN structure. SAFNet [46] introduced Adaptive Feature Fusion and Self-Enhanced Modules. However, some problems remain in the FPN structure. These methods all use sequential local fusion, which is an indirect interaction method and ignores the loss in the information flow process. Inspired by [14], all feature map layers present meaningful information for object detection. Therefore, designing a direct interaction and global fusion structure is of great significance for multi-scale object detection.
III Methods
This paper introduces SAR-Net, a SAR object detection framework based on a unity compensation mechanism. Built upon the YOLOv6 [47] architecture, SAR-Net consists of three components: backbone, neck, and head. Leveraging the advanced detection head of YOLOv6, we retain its head in the SAR-Net framework while focusing on innovations in the backbone and neck. In this section, we first present the overall architecture of SAR-Net, followed by an introduction to our proposed unity compensation mechanism, which addresses the issue of hindered information interaction due to local information fusion. Finally, we clarify the newly designed direction-aware attention module embedded in the backbone and neck, dedicated to acquiring object directional and positional information.
Figure 3: Overview of backbone components. DAM represents the Direction-aware Attention Module.
III-A Overall Framework of SAR-Net
The overall framework of SAR-Net is illustrated in Fig. 2, where the backbone and neck components are specifically tailored for SAR images. Specifically, considering that SAR image content is predominantly background, and detected objects are directional, using feature extraction networks designed for optical images directly cannot capture the directional and positional information of objects under complex background interference. Therefore, we have designed a direction-aware feature extraction backbone network, as shown in Fig. 3. It embeds our designed direction-aware attention module, reducing background interference while capturing fine-grained information, and providing effective assistance for the subsequent global fusion of multi-scale features in SAR images. Additionally, to better address the multi-scale challenge in SAR images and avoid information dilution caused by traditional FPN local fusion, we propose a unity compensation mechanism. This mechanism comprehensively fuses deep and shallow semantic features at different scales, enabling each layer to focus on global information and providing robust prior knowledge for the detection head. The unity compensation mechanism will be detailed in the next section.
III-B Unity Compensation Mechanism
In our implementation, the process of unity and compensation corresponds to three modules: the multi-scale alignment module, the multi-level fusion module, and the multi-feature embedding module. These modules collaborate to unitedly compensate for features, we illustrate the detailed information in the following subsections. As shown in Fig. 2, the input to the neck includes the feature maps $C2$, $C3$, $C4$, and $C5$ extracted from the backbone, where $Ci\in\mathbb{R}^{N\times\tilde{C_{Ci}}\times H_{Ci}\times W_{Ci}}$. The batch size is denoted by $N$, the channels by $C$, and the sizes by $H\times W$. Additionally, the feature map sizes decrease progressively from $C2$ to $C5$, while the channel numbers double at each level.
Figure 4: Multi-level Fusion Module at both shallow and deep levels, here, MHA and FFN denote Multi-Head Attention and Feed-Forward Network, respectively, with N of them combined in transformer blocks. The multi-scale alignment process is also depicted on the left side of the diagram, and the concatenation operation is performed along the channel dimension.
III-B1 Multi-scale Alignment Module
Given the crucial premise that both shallow texture and deep semantic information of objects in SAR images are important, we choose the outputs $C2$, $C3$, $C4$, and $C5$ from the backbone for global feature alignment. We take two factors into consideration for aligning feature size: the computational complexity of global fusion and the necessity of preserving the fine-grained information extracted by the backbone. The former tends to favor smaller feature maps to reduce computational load, while the latter prefers larger feature maps to preserve detailed information. To balance these two aspects, the second smallest feature size is chosen as aligned size when four features need to be aligned while the smallest feature size is selected when aligning three features. We employ average pooling and bilinear interpolation operations to adjust the feature sizes to the aligned feature size, obtaining $\mathcal{F}_{global\_align}$. This module ensures effective aggregation of information through global alignment while preserving high-resolution features of small objects. The formula is as follows:
$\mathcal{F}_{global\_align}=[AvgPool(Xi),Xj,Interpolate(Xk)]_{c},$ (1)
here, $\left[...\right]_{c}$ denotes the concatenation operation along the channel dimensions. $Xi$, $Xj$, and $Xk$ represent the different multi-scale features to be aligned at the same stage.
III-B2 Multi-level Fusion Module
Due to the typically high resolution and rich detailed information in SAR images, small objects manifest as shallow texture details, while large objects exhibit deep semantic information. Performing fusion only once cannot fully preserve the information of objects at different scales. Therefore, to enhance the model’s ability to detect multi-scale objects, we designed two multi-level fusion modules: the Shallow Multi-level Fusion Module (Shallow-MFM) and the Deep Multi-level Fusion Module (Deep-MFM). Shallow-MFM. Given that convolutional operations can effectively extract local features and texture information, which is crucial for detecting small objects in SAR images, Shallow-MFM utilizes Direction-aware Attention Block (DA Block). As shown in Fig. 4 (a), these blocks ensure the comprehensive capture and fusion of global detailed information in SAR images, followed by a separation operation. The formula is as follows:
$\mathcal{F}_{global\_fuse}=DA\;Block(\mathcal{F}_{global\_align}),$ (2)
$\mathcal{F}_{emb\_K3},\mathcal{F}_{emb\_K4}=Seperate(\mathcal{F}_{global\_fuse}).$ (3)
As shown in Fig. 2, for better control of the information propagation and fusion process, we only separate two features and compensate them onto the input features $C3$ and $C4$. This is because $C5$ contains the position and feature information of large objects, while the compensated features encompass more information about small objects. Simultaneously, the feature information from $C2$ has already been sufficiently utilized, and we no longer use it in the subsequent fusion process. Deep-MFM. Due to the advantages of transformers in handling global relationships and long-distance dependencies, and considering that large objects require more global information for accurate detection and localization, we employ a transformer fusion module in Deep-MFM. As illustrated in Fig. 4 (b), it consists of a stack of $N$ transformer blocks, where each transformer block includes Multi-Head Attention (MHA) and a Feed-Forward Network (FFN). We follow the configuration for Multi-Head Attention from [48], and this part of the process is expressed by the formula:
$\mathcal{F}_{global\_fuse}=Transformer(\mathcal{F}_{global\_align}),$ (4)
$\mathcal{F}_{emb\_B3},\mathcal{F}_{emb\_B4}=Seperate(\mathcal{F}_{global\_fuse}).$ (5)
Simultaneously, in constructing the FFN, we utilize depth-wise convolution, which can be considered as an implicit channel attention mechanism, enhancing the local connections of the transformer block. It is represented as follows:
$\mathcal{F}_{FFN}=CB_{2}(\psi(DW_{conv}(CB_{1}(X)))),$ (6)
here, $CB_{i}$ denotes the $i$-th convolution and batch normalization operation, $DW_{conv}$ represents Depth-Wise convolution, and $\psi$ is the ReLU activation function. Before merging the output features, to retain information about small objects in $K3$, we compensate the separated features onto the intermediate features $K4$ and $K5$, providing the detection head with output features rich in information about both small and large objects.
Figure 5: Architecture of Multi-feature Embedding Module, where Resize represents the average pooling operation or bilinear interpolation.
Figure 6: Architecture of Direction-aware Attention Module. The concatenation operation is performed along the spatial dimension.
III-B3 Multi-feature Embedding Module
Inspired by the fusion architecture with two branches in [49], to embed global information more effectively at different levels, we employ attention operations for information fusion, as illustrated in Fig. 5 Specifically, we input local information (referring to the features at the current level) and globally embedded information (generated by MFM) into two separate 1D convolutions for feature transformation. The Sigmoid function is utilized to generate attention weights from global information, which are then applied to weight the local information. Additionally, for further compensation, a residual connection is employed. This is aimed at achieving deep fusion and enhancement of global and local information. Finally, the DA Block is used for further feature extraction and fusion.
III-C Direction-aware Attention Module
Research on object detection in SAR images has shown that attention mechanisms can significantly enhance model performance. However, existing attention mechanisms often concentrate solely on positional information, frequently neglecting directional cues. This results in a lack of a method that seamlessly integrates both directional and positional information and embeds them into the attention map. In response to this, we propose the direction-aware attention module. As shown in Fig. 6, it can be regarded as a means of enhancing the feature map’s representational capacity, obtaining direction-aware feature maps through two steps: Direction-aware Generation and Channel Attention Embedding. These steps contribute to facilitating the subsequent extraction of object features. Below, we provide a detailed description of these two steps.
III-C1 Direction-aware Generation
Due to the dominance of background in SAR image content and the presence of interference areas such as docks, airports, and terminals, directly locating and detecting objects becomes challenging. Additionally, SAR images often exhibit the edge contours of objects, reflecting distinct directional features, while extracting detailed information within the objects is challenging. Therefore, leveraging the stability and distinctiveness of object direction information can effectively reduce the difficulty of localization. Based on the analysis above, deformable convolution, capable of adapting to and describing object deformations well, is employed. We perform deformable convolution along one direction (either the row or column direction) on the input features, generating two directional attention maps. These attention maps are then used to modulate the input features, and this process can be expressed as:
$X_{direction}=(\sigma(DConv_{row}(X)+DConv_{column}(X)))*X,$ (7)
here, $X$ represents the input features, and $\sigma$ denotes the Sigmoid function. This bidirectional convolution attention mechanism facilitates the extraction of directional information from low-level features in two directions, thereby guiding the refinement of input features. To further leverage the directional attention maps and avoid capturing only local contextual information, we forego using large-kernel convolutions and instead employ two parallel $1D$ feature encoding approaches. These methods independently aggregate input features in the vertical and horizontal directions, resulting in two separate localization-aware feature maps, as illustrated below:
$Z^{h}=\frac{1}{W}\sum_{i=0}^{W}X(h,i)\quad h=0,1,...,H,$ (8)
$Z^{w}=\frac{1}{H}\sum_{j=0}^{H}X(j,w)\quad w=0,1,...,W,$ (9)
the output feature maps $Z^{h}$ and $Z^{w}$ encapsulate global statistical information along the height and width directions, effectively aggregating multi-directional global contextual features of the input. This fusion of spatial positional information proves beneficial for subsequent detection tasks.
III-C2 Channel Attention Embedding
The aforementioned process enables a global receptive field and encodes precise directional awareness. To fully leverage this information and capture inter-channel relationships, we first align the two feature maps along the spatial dimensions. Subsequently, they are fed into the convolutional transformation function CBR, yielding:
$g=CBR([Z^{h},Z^{w}]_{s}),$ (10)
where $\left[...\right]_{s}$ denotes the concatenation operation along the spatial dimensions, and CBR represents convolution, batch normalization, and rectified linear unit (RELU) activation function. It encodes the features with both horizontal and vertical directional information after reintegration. Subsequently, we decompose the concatenated features into $g^{h}$ and $g^{w}$ along the spatial dimension. Two $1D$ convolutions are employed to restore the feature dimensions, and finally, the weight information from the Sigmoid function is applied to the input features, resulting in:
$Y=X*\sigma\left(Conv_{h}(g^{h})\right)*\sigma(Conv_{w}(g^{w})).$ (11)
In SAR object detection, we leverage the directional nature of SAR image objects by applying attention mechanisms to input features from two directions, extracting features in both horizontal and vertical directions. Simultaneously, we use average pooling and convolution to reduce the spatial scale and channel dimensions of intermediate features, reducing computational complexity in the embedding process. Compared to traditional methods, this encoding process enables our directional attention to more accurately identify objects of interest and achieve more efficient object detection.
III-D Loss Function
The overall loss of our network consists of two components [47]: the loss of object classification $L_{cls}$ and the loss related to bounding box regression $L_{reg}$. It can be expressed as:
$L_{total}=L_{cls}+L_{reg},$ (12)
we adopt the VariFocal Loss (VFL) [50] as the classification loss $L_{cls}$. VFL balances the learning signals between positive and negative samples by considering their varying levels of importance. The bounding box regression loss $L_{reg}$ comprises IOU loss and probability loss, as shown below:
$L_{reg}=\lambda_{iou}L_{iou}+\lambda_{dfl}L_{dfl},$ (13)
here, $\lambda_{iou}$ and $\lambda_{dfl}$ are hyperparameters. We choose the advanced SIOU [51] as the IOU loss $L_{iou}$. Due to the consideration of ambiguity and uncertainty in the data without introducing additional strong priors, the Distribution Focal Loss (DFL) [52] contributes to improving box localization accuracy. Therefore, we introduce an additional loss $L_{dfl}$, which is attributed to the bounding box regression loss.
IV Experiments and Analysis
To verify the superiority of SAR-Net, we perform extensive experiments on three public benchmark SAR object detection datasets, a fine-grained aircraft detection dataset SAR-AIRcraft-1.0 [53] and two ship detection datasets SSDD [54], HRSID [55]. Furthermore, we compared our proposed SAR-Net with the existing SOTA detection methods, and the performance of SAR-Net proved to be superior.
TABLE I: The Number of Instance Objects of Different Categories on the SAR-AIRcraft-1.0
Categories Train Test Total
A330 278 31 309
A320/321 1719 52 1771
A220 3270 460 3730
ARJ21 825 362 1187
Boeing737 2007 550 2557
Boeing787 2191 454 2645
other 3223 1041 4264
Total 13513 2950 16463
Figure 7: Bounding box sizes distribution of three SAR Datasets, where (a), (b), and (c) represent SAR-AIRcraft-1.0, SSDD, and HRSID datasets respectively.
IV-A Dataset and Implementation Details
To most accurately evaluate the performance and generalizability of our SAR-Net framework, we conducted extensive validation experiments on a recently released fine-grained multi-class SAR aircraft object detection dataset and two commonly used SAR ship object detection datasets.
TABLE II: Ablation Study of Three Modules on Unity Compensation Mechanism
[…]
86.5 0.813 4.64 17.80
$\checkmark$ 87.1 0.829 4.18 16.51
$\checkmark$ 85.9 0.819 4.69 15.78
$\checkmark$ 86.8 0.827 4.80 17.53
$\checkmark$ $\checkmark$ 87.8 0.835 4.87 16.82
$\checkmark$ $\checkmark$ $\checkmark$ 88.5 0.846 5.67 18.84
IV-A1 SAR-AIRcraft-1.0
All images in the SAR-AIRcraft-1.0 dataset are collected from the Gaofen-3 satellite. The polarization method is single polarization, the spatial resolution is $1m$, and the imaging mode is the spotlight. Considering the size of the airport and the number of parked aircraft, the dataset mainly uses image data of three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport, including a total of $800\times 800$, $1000\times 1000$, $1200\times 1200$ and $1500\times 1500$. 4 different sizes, a total of 4,368 pictures and 16,463 aircraft object instances. The specific categories of aircraft include A220, A320/321, A330, ARJ21, Boeing737, Boeing787, other. The number of instance objects of different categories are shown in Table I, where other represents aircraft instances that do not belong to the other 6 categories. The size distribution of the aircraft object slices in the dataset spans a wide range. As shown in Fig. 7 (a), some objects are smaller than $50\times 50$, and some aircraft objects are larger than $100\times 100$. The overall object size is multi-scale. In addition, in this paper, we use the default dataset partitioning by [53] and resize all images to $800\times 800$ pixels. For the sake of fairness, we adopt the above settings when compared with other methods.
TABLE III: Research on the Effectiveness and Optimal Embedding Locations of Direction-aware Attention Module
[…]
Baseline 86.5 0.813 4.64 17.80
+Row Attention 87.2 0.825 4.89 18.83
+Column Attention 86.9 0.824 4.89 18.83
Behind RELU 87.4 0.831 4.82 18.69
Behind ERBlock 87.1 0.819 4.79 18.35
DAM 87.8 0.834 5.06 19.76
IV-A2 SSDD
The SSDD dataset comprises data primarily sourced from Radarsat-2, TerraSAR-X, and Sentinel-1 sensors, featuring four polarization modes: HH, HV, VV, and VH. SAR image resolutions range from $1m$ to $15m$. Comprising 1,160 SAR images and 2,456 ship objects, each image contains an average of 2.12 ship objects. The height-width distribution of each ship object is illustrated in Fig. 7 (b). For standardized benchmarking, the SAR images with serial numbers ending in 1 and 9 in the SSDD dataset form the test set, ensuring an 8:2 ratio division with the training set. Consequently, the training set consists of 928 images, and the test set comprises 232 images. To facilitate fair comparisons with other methods, all images are resized to $640\times 640$ pixels in all experiments, maintaining consistent data distribution. This approach ensures a level playing field for evaluating the performance of different methods on the SSDD dataset.
TABLE IV: Collaborative Research Within the Framework, UC Represents Unity Compensation
[…]
86.5 0.813 4.64 17.80
88.5 0.846 5.67 18.84
87.8 0.834 5.06 19.76
90.2 0.856 6.09 20.80
IV-A3 HRSID
To assess the ship detection capabilities of our proposed method across different datasets, we conducted performance validation on the HRSID dataset—a large-scale, high-resolution SAR image ship object detection dataset designed for tasks such as object detection and instance segmentation. The HRSID dataset encompasses SAR images with diverse properties, including variations in sea area, polarization, sea condition, resolution, and coastal port scenarios. As depicted in Fig. 7 (c), the dataset encompasses ships at multiple scales, annotated with bounding boxes within various environments (e.g., different polarization modes, sensor types, and scenes). Panoramic images with resolutions ranging from $1m$ to $5m$ were cropped to a uniform size of $800\times 800$. In total, the HRSID dataset consists of 5,604 cropped SAR images, containing 16,951 ships. These images are distributed in a ratio of 6.5:3.5 between the training and test sets. This dataset diversity ensures comprehensive evaluation and validation of our proposed method’s performance under various conditions and scenarios. Our proposed framework is implemented in PyTorch and runs on a workstation with an NVIDIA A100 GPU. During training, the batch size was set to 32, and the initial learning rate was 0.02, reduced to 0.0002 in a cosine decay manner, and trained for a total of 300 epochs to ensure it fully converged. Standard Stochastic Gradient Descent (SGD) was used as the optimizer, and the parameters were all at default values. For input images during training, we only used mosaics augment, discarding Hue Saturation Values (HSV) and random flipping as they are not beneficial for SAR images.
TABLE V: Comparison with Different FPN Variants
[…]
FPN[11] 87.4 0.833 4.55 17.60
PAFPN[56] 87.2 0.828 4.82 18.13
AFPN[43] 85.2 0.807 4.67 18.19
RepBi-PAN[47] 86.5 0.813 4.64 17.80
UC Mechanism 88.5 0.846 5.67 18.84
IV-B Accuracy Metrics
The accuracy assessment involves evaluating the concordance and distinctions between the detection outcome and the reference label. Key accuracy metrics such as recall, precision, and mAP (mean Average Precision) are employed to assess the performance of the compared methods. The computations for precision and recall metrics are delineated as follows:
$Precision=\frac{TP}{TP+FP},$ (14)
$Recall=\frac{TP}{TP+FN},$ (15)
in the equations, the terms true positive (TP) and true negative (TN) represent accurate predictions, while false positive (FP) and false negative (FN) denote incorrect outcomes. Precision and recall are associated with commission and omission errors, respectively. The F1-score assesses both precision and recall rates, and it can be expressed as:
$F1\textendash score=2\times\frac{Precision\cdot Recall}{Precision+Recall}.$ (16)
The mAP serves as a holistic indicator derived by averaging the values of Average Precision (AP) and utilizes an integral approach to compute the area under the Precision-Recall curve across all categories. The mAP calculation can be expressed as:
[…]
in the formula, $p$ represents Precision, $r$ represents Recall, and $N$ stands for the number of categories. In addition, model parameter sizes and GFLOPs are also important indicators used to evaluate the model, which are related to the implementation and practical application of the model.
TABLE VI: Comparison with Other Methods on SAR-AIRcraft-1.0. The Accuracy of Each Class is Represented by mAP with an Intersection Over the Union Threshold of 0.5
[…]
Faster R-CNN[57] 77.3 97.0 86.4 74.0 59.5 74.0 71.2 77.2 0.746 41.09 142.73
Cascade R-CNN[58] 75.2 97.7 88.6 79.5 55.0 69.6 69.8 76.9 0.728 69.11 185.89
YOLOv6n[47] 92.6 97.5 96.2 81.4 78.3 80.6 78.9 86.5 0.813 4.64 17.80
YOLOv8n[59] 90.2 97.8 96.5 84.3 81.7 79.9 83.5 87.7 0.830 3.15 13.57
SKG-Net[60] 66.0 79.9 79.2 66.1 64.8 70.4 72.1 71.2 0.659 31.38 103.69
SA-Net[53] 80.3 94.3 88.6 78.6 59.7 70.8 71.3 77.7 - - -
MLSDNet[61] 85.1 96.9 91.5 83.2 71.7 72.1 78.4 82.7 0.779 1.42 6.72
HRLE-Det[62] 90.6 98.1 98.1 82.6 78.9 79.1 82.5 87.1 0.840 3.02 14.20
SAR-Net 91.0 98.5 96.9 92.6 86.6 82.4 83.6 90.2 0.856 6.09 20.80
TABLE VII: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on SSDD
[…]
Faster R-CNN[57] 96.5 0.953 41.09 91.33
Cascade R-CNN[58] 97.0 0.955 69.11 118.90
YOLOv6n[47] 97.9 0.949 4.64 11.40
YOLOv8n[59] 98.1 0.956 3.15 8.70
LPEDet[63] 97.5 0.938 5.66 18.34
MLSDNet[61] 97.2 0.937 1.42 4.31
HRLE-SARDet[62] 97.6 0.948 3.02 9.10
YOLOv8m[59] 98.8 0.968 25.40 78.92
SAR-Net 98.7 0.971 6.09 13.76
IV-C Ablation Study
To comprehensively and detailedly analyze the effectiveness of SAR-Net, we conducted extensive ablation experiments on the SAR-AIRcraft-1.0 dataset with different framework designs and internal modules. The best results are highlighted in bold.
IV-C1 Impact of Unity Compensation Mechanism
[…]
TABLE VIII: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on HRSID
[…]
Faster R-CNN[57] 81.0 0.742 41.09 142.73
Cascade R-CNN[58] 82.7 0.813 69.11 185.89
YOLOv6n[47] 91.2 0.873 4.64 17.80
YOLOv8n[59] 91.4 0.877 3.15 13.57
LPEDet[63] 89.9 0.862 5.66 28.61
MLSDNet[61] 89.1 0.845 1.42 6.72
HRLE-SARDet[62] 90.8 0.866 3.02 14.20
CMFT[64] 89.6 0.859 - -
YOLOv8m[59] 93.0 0.896 25.40 123.12
SAR-Net 92.8 0.898 6.09 20.80
IV-C2 Impact of Direction-aware Attention Module
To prove the effectiveness of the proposed direction-aware attention module, we conducted a series of detailed experiments, as shown in Table III. We split the attention modules in two directions and embedded them in the YOLOv6n backbone separately, and both received performance improvements. In addition, we tried to embed DAM in different positions of the backbone. Except for the embedding position shown in Fig. 3, we also tried to move the DAM behind the RELU and each ERBlock, and obtained 0.9% and 0.6% performance improvement respectively with a small increase in parameters. Exhaustive experiments allowed us to determine the most suitable embedding position. These experimental results show that the acquisition of direction-aware information is significantly helpful for SAR images containing directional objects.
IV-C3 Collaboration of the Framework
The effectiveness of the components that make up SAR-Net has been demonstrated, and to further explore the synergy between them, we combined them and found the desired results. For the unity compensation mechanism, the backbone that can extract direction-aware information gives full play to the advantages of global fusion and enriches the information of multi-scale features. For the DAM-embedded backbone, the unity compensation mechanism retains detailed information and enhances the efficiency of information fusion and transmission. As shown in Table IV, after combining the two, their collaborative complementarity is fully utilized and performance is further improved by 0.4%.
Figure 8: Visualization results of object detection under different methods on SAR-AIRcraft-1.0, (a)-(d) depict four randomly selected images, where differently colored rectangles represent aircraft of various types. To enhance clarity, we have boldened the object boxes. Yellow ellipses indicate false positives, while red ellipses represent instances of false negatives and misclassifications.
IV-C4 Comparison of Different FPN Architecture
We chose the classic FPN structure and its latest variants AFPN [43] and RepBi-PAN [47] for comparison. We conducted the experiments based on YOLOv6n and the results are shown in Table V. There is a large accuracy gap when the parameters are similar. As we discussed in the section I, sequential local fusion will lead to information loss, especially for SAR images with large object scale variants. Top-down and bottom-up structures will lose a lot of detailed information and semantic information before detection. Global information fusion and compensation will retain the multi-scale information of the object to the greatest extent, therefore unity compensation mechanism achieves the best performance.
IV-D Comparisons with Previous Methods
To demonstrate the effectiveness and robustness of our proposed method, in addition to comparing it with classic and latest object detection methods on three datasets (SAR-AIRcraft-1.0, SSDD, HRSID), we also selected methods specially designed for SAR images. With the exception of the unreleased and non-detailed network structure of the comparative method, we successfully reproduced all other comparative methods. All methods are compared and evaluated from four aspects: mean Average Precision, F1-score, parameter size, and GFLOPs. The experimental parameter settings for each dataset remain consistent.
Figure 9: Visualization results of object detection under SAR-Net on SSDD.
Figure 10: Visualization results of object detection under SAR-Net on HRSID.
IV-D1 Comparison Results on SAR-AIRcraft-1.0
We used the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [57], Cascade R-CNN [58], and one-stage detectors: YOLOv6 [47], YOLOv8 [59]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [60], SA-Net [53] and MLSDNet [61], and a formidable model, HRLE-SARDet[62]. Table VI shows the detection results under different methods. In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As shown in Table VI, it can be seen that due to the fusion of global information, SAR-Net achieves the highest performance in almost every category, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, with a significant increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some types of objects, such as ARJ21, due to its small size, the detailed features obtained are not sufficient, resulting in low detection accuracy.
IV-D2 Comparison Results on SSDD
To verify the superiority of our method on different datasets, we selected the widely used SSDD dataset and compared it with some of the latest object detection methods and SAR ship detection methods. A variety of experiments ensured the rigor of our work. As indicated in the data presented in Table VII, our method exhibits strong competitiveness even when transferred to ship detection tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, our approach excels in information fusion and transmission, making it less prone to overfitting on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we experience only a marginal 0.1% decline in accuracy, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
IV-D3 Comparison Results on HRSID
To further validate the generalization and robustness of our method, we opted for the more challenging ship detection dataset, HRSID, in comparison to SSDD. HRSID presents greater difficulty with a higher number of instances, larger image resolutions, and significant variations in object scales. This dataset serves as a more stringent test of the model’s capabilities in feature extraction and utilization. Continuing our evaluation, we chose state-of-the-art object detection methods, including those specifically tailored for ship detection in SAR images. Compared to other approaches, our method maintains excellent performance with a relatively small computational cost. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits slightly lower accuracy but significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
IV-E Result Visualization
Due to the SAR-AIRcraft-1.0 dataset’s nature of fine-grained object detection, it presents certain detection challenges, leading to notable variations in detection results among different methods. To provide a more intuitive representation of the performance of various methods on SAR-AIRcraft-1.0, we selected a few images and visualized the detection results using different methods, as shown in Fig. 8. In comparison with other methods, our approach consistently identifies objects. While occasional false positives may occur, the incorporation of sufficient directional awareness information and the utilization of global fusion to minimize information loss result in minimal instances of false negatives in our detection results. For instance, the green bounding box (representing A220) in (b), the brown bounding box (representing ARJ21) in (c), and the orange bounding box (representing Boeing787) in (d) were undetected by other methods, highlighting the superiority of our approach in these scenarios. For the SSDD and HRSID datasets, given their high baseline levels of average precision and the single-category nature of object detection, we provide visualization results exclusively for SAR-Net. The visualizations for SSDD and HRSID are presented in Fig. 9 and Fig. 10, respectively. SAR-Net demonstrates excellent detection performance in both inshore and offshore backgrounds.
V Conclusion
As the focus on deep learning-based remote sensing tasks grows, SAR image object detection has become a crucial research direction. Establishing a complementary relationship between high-level semantic information and low-level detailed information without losing critical details is essential for solving the multi-scale object feature fusion problem. In response to this, we propose a framework capable of globally fusing multi-scale direction-aware information, named SAR-Net. This framework effectively addresses the challenges of information loss due to sequential local fusion and the difficulty in capturing detailed information under complex background interference, allowing multi-scale detection features to focus on global information. Experiments conducted on SAR image aircraft and ship datasets demonstrate the effectiveness and excellent generalization of our framework. We believe this global information fusion perspective will provide valuable insights for SAR image detection tasks.